OpenStudy (anonymous):

8(2x – 14) + 13 = 4x – 27 i know x=6 but i need a step by step way how to get it

OpenStudy (freckles):

I would distribute where possible 8(2x-14)=? recall a(b+c)=a*b+a*c

OpenStudy (anonymous):

16x-112

OpenStudy (freckles):

so then we have: \[16x-112+13=4x-27\]

OpenStudy (freckles):

now we do have like terms on the left hand side

OpenStudy (freckles):

-112+13=?

OpenStudy (anonymous):

-99

OpenStudy (freckles):

that is what I have too... \[16x-99=4x-27\]

OpenStudy (freckles):

now we need to get the terms with x on one side and the terms without x on the opposing side

OpenStudy (anonymous):

-4x?

OpenStudy (freckles):

right subtract 4x on both sides

OpenStudy (freckles):

16x-4x-99=4x-4x-27 this will make x go away on one side since 4x-4x=0 12x-99=-27

OpenStudy (freckles):

you mean +99 on both side sright?

OpenStudy (anonymous):

haha yea

OpenStudy (freckles):

\[12x-99+99=-27+99\] yes this would isolate the 12x on the left hand side since -99+99=0 \[12x=-27+99\]

OpenStudy (freckles):

so we do have like terms on the right hand side -27+99=?

OpenStudy (anonymous):

combine like terms

OpenStudy (freckles):

-27+99=72 so we have: \[12x=72 \text{ since } 99-27=72\] last step is to...

OpenStudy (freckles):

we want to isolate the x but x is being multiplied by 12

OpenStudy (freckles):

so divide both sides by?

OpenStudy (anonymous):

12

OpenStudy (freckles):

\[12x=72 \\ \frac{12x}{12}=\frac{72}{12} \\ \frac{12}{12}x=\frac{72}{12} \\ 1 x=\frac{72}{12} \text{ since } \frac{12}{12}=1 \\ x=\frac{72}{12} \text{ since } 1x=x\]

OpenStudy (anonymous):

thanks so much

OpenStudy (freckles):

you just have to reduce 72/12 and you are done (aka divide)

OpenStudy (freckles):

np